Oxford Materials IT Committee Meeting

MICHAELMAS TERM 2014
A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 28th October 2014 in the Hume-Rothery Meeting Room.

AGENDA

  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters outstanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting: Monday 2nd February 2015 at 11am in ETB Wolfson meeting room.

ITEM 2. MATTERS OUTSTANDING FROM THE PREVIOUS MEETING

See minutes below.

ITEM 3: MATTERS FOR DISCUSSION

Verbal reports on the status of various projects:
IT services projects: Digital strategy, RDM, SITS/BIZTALK, WDE, TONE, maillists, Chorus, ISP
Departmental projects: RexRichards, EM, Phones, Photocopiers, iLab
Information from other Committees: MPLS ICT Panel, Divisional Software Licensing Group

ITEM 4: HEAD OF IT REPORT

The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: Khalid Schofield has been away on unpaid leave since 8th August and there is still no sign of him returning soon. Alex continues to provide some support remotely.

Admin Network Emma Purves replaced Nicky Carlin.

Teaching Laboratory The very old laptops have finally been replaced with six new laptops (plus a spare). These laptops have already been used with the new Matlab/Labview practical.

A labelling/testing error during electrical rewiring works in July resulted in numerous devices being fried by 415V, causing considerable extra work for IT staff. The following IT devices mostly in the teaching laboratory were replaced (under insurance): 6 optical microscope control computers; 4 computer room desktops; 7 monitors; 6-way VGA switcher; 48port gigabit network switch; storage array UPS.

Lecture theatres / meeting rooms The HRMR computer was replaced from trickledown stock and the polycom communicator replaced with a Chat160 audio unit.

Digital Print-room The two computers were replaced from trickledown stock. One of the spare A4 colour printers was almost immediately used to repair a major failure on the Rex Richards printer.

Library The self-issue system still needs upgrading. Annual Part2 thesis rebinding done.

MML All desktops are now running Ubuntu (rather than OpenSUSE. Rothery2 server is ready to decommission. Landau was reinstalled. Zener has been closed to users but still awaits reinstallation. Rex Richards Aircon broke again and was fixed yet again – minimal downtime. Kittel has been rather unstable over the summer requiring rebooting on several occasions and several more nodes have developed hardware problems (out of warranty).

FIM Phil Edmondson moved to USA. New storage array Onkalo. LEAP now running Windows7. Five new desktops and associated rearrangement of systems.

EM Adam Papworth moved to Gatan. Laptops ordered for EM staff. Over the summer considerable effort has gone into the migration to the new iLab equipment booking system which went live on 27th October. User documentation was written covering the booking system and how to use the private network infrastructure. The Begbroke private network still needs further refinement when time permits.

MFFP Four new HPC systems purchased. Further hard disks failed and were replaced in the TJM storage array (wrong type of disks in first unit!). Numerous security updates to the DSM storage control software have been applied over the summer attempting to coincide with patch Tuesdays in order to minimize downtime of the array which now provides storage to 13 HPC systems and numerous user desktops (this has become a bit of a management nightmare to liaise with so many users in order to avoid data loss). The old storage array is now full (90+Tb) so TJM group have just purchased a new array and AJW group are also purchasing a new array. These storage array(s) and the HPC systems are all due to be relocated to the Hume-Rothery server room.

OMCS OMCS equipment computers are still being moved behind the private network. Jenni Crosskey replaced Laura Restorick. Peter Northover retired.

Rex Richards / Peter Bruce: refurbishment of the ground floor and basement of Rex Richards is completing this week and Professor Bruce’s group of 25 people are due to start arriving next week. The small meeting room AV equipment specification is still not quite finalized. Items requiring attention include: Salto door entry system, wireless, networking, phones, printing, photocopier, network registrations.

Website www-solar.materials was fixed and updated. All CMS sites were upgraded to 1.11.11. (CMSMS 2.0 has been released for beta testing). SSL certificates need replacing. Also security updates to drupal mffp.materials website. FNG group split into NSM group (JHW) and CNano groups (KP) so the websites were rearranged accordingly.

Publications X sets of business cards and several visa passport photographs. Mugshot display boards are still being updated for the new academic year. Printed 60 copies of graduate handbook, 50 copies of Prelims and Finals handbooks. Awaiting Department Handbook. Weekly editorial of Departmental Newsletter. Signage and posterboards for two workshops (JRY, SGR) and opendays. Framing of photographs for EM opening ceremony.

Network Network switches have (mostly) been migrated to a private subnet for security and management, releasing 30+ public IP addresses which are in short supply as the department continues to expand (but our IP allocation remains limited to 1024 addresses). Networking in PR was all replaced with gigabit switching. The newly refurbished Rex Richards ground floor area network is now live and will be populated when users arrive in November.

Security Oxcert have released more than 100 security bulletins since the last meeting. Several of these were major problems (e.g. Heartbleed, Shellshock, Poodle) and required IT staff to check and patch hundreds of systems. There have been no known compromises. IT staff have dealt with a handful of users with browser hijacks by malware.

WDE: WDE pilot project ended in June so Jayne and Philippa were migrated to the new WDE service. IT staff have received WDE training. So far only six user’ devices encrypted – need to expand uptake targeting senior staff.

Disposal The old post-office storage room outside 21BR has been cleared out and shelving installed so that it can act as a store room for WEEE equipment awaiting disposal. The disposal scheduled for September did not happen due to lack of man power.

Software PSG group purchased another floating license for COMSOL+RFmodule. Matlab is now fully site-licensed. A new floating license for Intel Fortran Compiler for Windows was purchased. All users of the old license need to migrate to the new license server so that rothery2 can be decommissioned. End of Redhat site license so migrated oums-sfgpu and oums-ocampo to CentOS.

Disaster Recovery and Backup Numerous disk fails have occurred and data has been restored from backups where possible. A couple of systems were found to no-longer have valid backup.

Documentation updates Documentation on the private Mediawiki continues to expand, but time for documentation has been limited due to being short staffed.

New Device registrations May 16; June 5; July 11; August 17; September 25; October 25


MINUTES

A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 3rd June 2014 in the Hume-Rothery Meeting Room. The following people attended: APD, PJW, CA, RT, JRY, SCB, LJ. Apologies: RS.

1. Apologies for Absence and Membership of the Committee

Khalid Schofield was out of the country. Rob Saunders was on half-term holiday.

Lewys Jones has joined the committee at this meeting as a postdoctoral representative.

2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) *PC maintenance – Equinox computer hardware breakdown service (Minute 2a)
A process is still required for (automatic?) notification of owners that their warranty is due to expire. OUCS no-longer offer this service however their supplier Equinox continues to offer cost effective insurance if warranty is required – such as for laptops.

Action: PJW

b) *Photocopying and Printing (Minute 2b)
All networked printers have been created in the new Print server and the drivers tested. Instructions for installing printers have been updated for the new print server. The old print server is almost ready to turn off. Users will be notified via the Newsletter and via email about a changeover day which may require users to reinstall any printers still on the old iPrint server.

Action: PJW

c) Draft Information Security Policy (Minute 2f)
Again there has been little opportunity to progress towards formalising the departmental Information Security Policy for approval by the ITC and subsequently DC. However PJW noted that the ISP documentation and toolkit has evolved considerably over the last few months and PJW is keen to revisit implementing a departmental IS policy before Trinity Term. (not progressed)

Action: PJW

The university now has an online information security training module available at http://www.it.ox.ac.uk/infosec/awareness/ which provides guidance to staff about best practice information handling.

d) XP end of life (Minute 2e) As of 15th October 2014 there are still 48 systems requiring attention (7 user desktops/laptops; 4 virtualmachines; 5 control systems; 42 equipment of which 26 are still on public network; 6 home devices no-longer registered for department network;)

Action: PJW

e) Office 2003 end of life planning (Minute 2f)
The number of computers running Office 2003 has decreased from 47 to 26 (8 to be reinstalled, 18 equipment to be upgraded to office 2010).

Action: PJW

f) *Fresher survey 2013 (Minute 2i)
The department should review (and document) current wireless coverage and generate a plan for increased coverage, also investigating possible alternative cheaper solutions for providing wireless (currently £600/AP).

A set of six cheaper access points (£200/AP) have been purchased and they will be piloted in the Hume-Rothery Building, releasing the four existing Hume-Rothery building AP’s for redeployment elsewhere around the Department. APD asked to be involved with identifying locations for increased wireless coverage.

Action: KS/RS

g) Lecture Theatres (Minute 6b)
The computer in Hume-Rothery meeting room was replaced from trickledown stock and the polycom communicator was replaced with a Chat160 resulting in much improved skype performance for interviews.

h) Copyright notices and illegal downloads (Minute 6c)
All users were reminded via the department newsletter that copyright infringement is against the law (as well as being against university rules) and using university facilities for such illegal downloading of copyright materials is detrimental to the reputation of the university and risks affecting our network access.

3 Matters for Discussion

a) Verbal update on various projects

PJW updated the committee verbally upon the status of the following projects:

  • IT services projects: Digital strategy, RDM, SITS/BIZTALK, WDE, TONE, maillists, Chorus, ISP
  • Departmental projects: RexRichards, EM, Phones, Photocopiers, iLab
  • Information from other Committees: MPLS ICT Panel, Divisional Software Licensing Group

4. Head of IT Report
A written report was circulated with the agenda.

5. *Finances
No information was presented at the meeting however Tim has supplied a full R12 transaction listing. Paul still needs to journal costs for stock computers from Aug/Sept/Oct.

6. AOB

a) Lecture theatres
Lecture theatres and meeting rooms all need technology upgrades. Also there is a desire for providing skype functionality in BR and ETB committee rooms.

Action: PJW / CA

7. Date of Next Meeting
TThe next meeting is on Monday 2nd February 2015 at 11am in ETB Wolfson meeting room.

The meeting closed at 12.15pm